OpenStudy (anonymous):

X=7y+2 in standard form?

OpenStudy (anonymous):

Standard form is \(ax + by = c\)

OpenStudy (anonymous):

x-7y=2

OpenStudy (anonymous):

You dont have to subtract 2?

OpenStudy (anonymous):

You don't have to subtract 2 if you subtract 7y instead. \[x = 7y + 2\]\[x - 7y = 2\]

OpenStudy (zale101):

subtract 7y from both sides

OpenStudy (anonymous):

Thank you, i have another one lol

OpenStudy (zale101):

post it :)

OpenStudy (anonymous):

1/3x-1/3y=-2

OpenStudy (zale101):

get rid of the fractions first

OpenStudy (zale101):

by multiplying 3 in both sides

OpenStudy (anonymous):

X-y=-6?

OpenStudy (zale101):

yeah

OpenStudy (zale101):

now put it on y=mx+b form

OpenStudy (anonymous):

Yaya ok want another one? Lol

OpenStudy (zale101):

meaning, add x on the other side y=x+6 to solve for y

OpenStudy (anonymous):

I dot have to do that just standard form it asks

OpenStudy (zale101):

we need to turn the equation to solve for y, that's what we are focusing at

OpenStudy (anonymous):

Dont* and oh okay, sorry im typing on my phone its a pain

OpenStudy (zale101):

that's fine

OpenStudy (anonymous):

0.25y=10c in standard form?

OpenStudy (zale101):

u mean 10c?

OpenStudy (anonymous):

Well actually it only says =10

OpenStudy (zale101):

if u written it this way, 0.25y=10 just divide both sides by 0.25 0.25/0.25=10/0.25

OpenStudy (zale101):

and you'll get y=40

OpenStudy (anonymous):

Ok thank youuuu :) got 2 more problems left lol

OpenStudy (zale101):

post em

OpenStudy (anonymous):

0.25x=0.1+0.2y?

OpenStudy (anonymous):

Y=4x-2?